OpenStudy (anonymous):

medal for the correct answer! Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

It's solve f= 1(x) ???

OpenStudy (anonymous):

Didn't get your question sorry

OpenStudy (anonymous):

@xoya it's solve f-1(x) when x = 4, so yes it would be solve f-1(x)

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

@xoya yeah, can you help me?

OpenStudy (anonymous):

I'm new to this question but I cant try :) Please wait and let me understand the question :)

OpenStudy (anonymous):

@sam801

OpenStudy (anonymous):

@xoya: what are the options 1 3 4 10 ?

OpenStudy (anonymous):

@sam801 the options are 1, 3, 4, and 10!

OpenStudy (anonymous):

the question is not clear for me, if it is like solve f - 1*x when x = 4 then it will be f(4) - 1*4 2*4 + 2 -4 8+2-4 which is equal to 6 but there is no option 6